A standing wave, generated on a string of length L, is represented by the following mathematical expression y(x,t) = 10 sin⁡(2πx) cos⁡(15πt). At what times would all the elements of the string have zero displacement [y(x,?) = 0]?

1
Expert's answer
2021-12-17T18:16:28-0500

"y(x,t) = 10\\sin\u2061 (2\u03c0x) \\cos\u2061(15\u03c0t),"

"y(x,t)=0,"

"\\cos\u2061(15\u03c0t)=0,"

"15\pi t=\frac{\pi}2+\pi n,~n\in \Z,"

"t=\frac1{30}+\frac n{15},~n\isin\Z."